Using the following atomic mass, Al= 27, S=32, O=16 what is the molar mass of aluminum sulfate?


1
Expert's answer
2021-10-27T06:44:46-0400

MM(Al2(SO4)3) "= 2 \times 27 + 3 \times 32 + 12 \times 16 = 54 + 96 +192 = 342 \;g/mol"